sun couriers, a parcel delivery company, has found that the delivery tome of parcels to Clients in the Durban metropolitan area after airport collection is normally distributed with a mean delivery time of 45 minutes and a standard deviation of 8 minutes. for a newly arrived consignment at Durban airport, what is the probability that a randomly selected parcel will take: between 45 and 51 minutes to deliver to the client.
1
Expert's answer
2019-04-23T12:22:44-0400

P(45<X<51)=P((45-45)/8<Z<(51-45)/8)=P(0<Z<0.75)=P(Z<0.75)-P(Z<0)=

=0.7734-0.5=0.2734.